# NeurIT: Pushing the Limit of Neural Inertial Tracking for Indoor Robotic IoT

### v1

@@ VI-A Overall Performance
ATE 16.75 26.04 1.72 1.76 1.66 3.94 1.63 1.40% ↑ 1.53 7.64% ↑
RTE 17.58 37.51 2.04 2.09 2.16 4.34 2.04 0.16% ↑ 2.02 1.19% ↑
PDE 0.2030 0.6836 0.0328 0.0319 0.0338 0.0688 0.0317 0.55% ↑ 0.0286 10.41% ↑
Seen AYE 127.97 114.15 51.39 57.52 70.95 63.74 27.21 47.05% ↑ 26.33 48.76% ↑
ATE 17.30 24.14 1.87 2.10 1.99 3.31 1.51 19.10% ↑ 1.51 19.10% ↑
RTE 18.02 40.67 1.87 2.18 1.92 3.20 1.66 11.16% ↑ 1.60 14.50% ↑
PDE 0.1799 0.7095 0.0291 0.0354 0.0310 0.0577 0.0278 4.62% ↑ 0.0243 16.55% ↑
RIDI Unseen AYE 130.59 105.93 48.47 53.67 67.42 64.64 24.34 49.78% ↑ 24.21 50.05% ↑
ATE 22.86 33.26 7.76 6.20 5.68 8.78 4.58 19.38% ↑ 4.37 23.10% ↑
RTE 30.49 35.10 3.19 3.17 2.77 6.43 2.60 6.03% ↑ 2.59 6.53% ↑
PDE 0.1240 0.1067 0.0432 0.0390 0.0267 0.0493 0.0216 18.99% ↑ 0.0181 31.99% ↑
Seen AYE 122.48 120.96 63.12 74.49 107.92 86.91 46.56 26.24% ↑ 47.60 24.59% ↑
ATE 17.30 33.47 8.91 7.73 6.38 10.27 5.78 9.52% ↑ 5.35 16.26% ↑
RTE 18.02 41.07 4.12 4.63 3.86 8.20 3.59 7.12% ↑ 3.55 8.05% ↑
PDE 0.0912 0.3166 0.0462 0.0358 0.0320 0.0592 0.0268 16.13% ↑ 0.0243 24.03% ↑
RONIN Unseen AYE 110.00 114.80 64.46 76.49 106.80 88.69 47.88 25.72% ↑ 46.66 27.61% ↑
Evaluation on RIDI & RONIN Datasets: While we mainly focus on robotic tracking in this paper, we also test NeurIT on pedestrian tracking datasets to verify the effectiveness of our model design. Tab. III shows the test results on RIDI and RONIN datasets. NeurIT achieves the best performance among all baseline models. While the baselines also achieve reasonable accuracy for seen data, their performance degrades significantly for unseen data, for which NeurIT exceeds them considerably. Specifically, in RIDI dataset, the performance gain increases from 17% on the seen test set to nearly 25% on the unseen test set. As for RONIN dataset, it’s worth noting that the authors only made half of the total datasets available. Consequently, the model might not be able to fully capture the essential features of pedestrians, leading to a slight decrease in performance gains on the test unseen conditions. The noticeably inferior results on the RONIN dataset, compared to the results on the RIDI dataset, can be attributed to this fact as well. In conclusion, the obvious enhancement achieved by NeurIT demonstrates the potential of NeurIT to serve as a unified model for both robotic and pedestrian tracking.
VI-B Benchmark Study
Performance over Time and Distance: In this study, we examine the growth of prediction errors over time and tracking distance. As depicted in Fig. 9a and Fig. 9b, our prediction errors, like other tracking systems, increase with time and distance. Nonetheless, NeurIT experiences a much slower growth rate compared with the best baseline approach (UniTS). For example, the ATE of NeurIT after tracking for 250 meters is still smaller than that of the best baseline at 50 m. At a distance of 250 meters or after 5 minutes of travel, NeurIT keeps the mean trajectory error to around 0.9m. In contrast, the best baseline model achieves a similar error in just 1 minute and reaches an error of approximately 2.6m after 5 minutes (around 250m). The results indicate the stability and robustness of NeurIT over long-time or long-distance tracking in indoor environments.
@@ VI-B Benchmark Study
Velocity Prediction Error: To assess NeurIT’s performance in speed estimation, we calculate the ground truth for each moment and evaluate the error in the predicted speed. As shown in Fig. 10a, NeurIT yields a consistently higher speed accuracy than the best baseline (UniTS), with a considerable margin of around 0.2 m/s. The error in speed estimation remains constant until the actual speed exceeds 1.2 m/s while remaining within a narrow range of less than 0.4 m/s. Insights from the velocity probability density function indicate that the error grows at higher speeds, potentially due to insufficient training data in that range. Augmenting the dataset with more high-speed instances could be a potential solution to address this issue. Conversely, the best baseline performs poorly in speed estimation, with an average error of over 2 m/s across all speed conditions. In Fig. 10b, a brief analysis of a tracking trajectory shows that NeurIT generates a precise and accurate path that closely matches the ground truth. In contrast, the trajectory produced by the top baseline model exhibits erratic and imprecise movement.
@@
Different Motion Status: The robot’s movement status typically adapts to environmental changes. To evaluate the robustness of NeurIT in various indoor settings, we performed additional experiments and visualized the predicted trajectories along with the corresponding speed in Fig. 11. The office in Building A, floor 1 is narrow and crowded, the robot can only follow the narrow corridors. During turning and straight-line movement, NeurIT demonstrates excellent tracking capabilities. It also exhibits reliable performance in challenging scenarios, including emergency starts, stops, and high-speed linear motion. Building A, floor 5 demonstrates an even more radical case with an extremely long and narrow corridor of approximately 40m. In the case of new environments, such as Building B and C, which represent complex and open areas, respectively, NeurIT is not affected by unseen conditions. These experiment outcomes showcase its robustness and adaptability in diverse indoor settings, even when confronted with previously unseen conditions.
VI-C Ablation Study
We proceed to evaluate the effectiveness of different design modules in NeurIT, focusing on four aspects: time-frequency input (TF), block-recurrent attention module (BR), multi-loss learning (m-loss), and a more detailed experiment on the impact of each loss function.
Multi-loss Learning: To evaluate the multi-loss learning, we first replace it with a single loss function, i.e., position loss as used in RONIN. The performance degrades to a certain extent across the four evaluation metrics. We observe some counter-intuitive phenomena from the test results. Adding the loss function $\mathcal{L}_{o}$ to $\mathcal{L}_{p}$ resulted in a deterioration of the model’s performance on ATE and RTE, despite an improvement in AYE from 35° to around 30°. This phenomenon is also observed with the loss combinations of $\mathcal{L}_{o}$ and $\mathcal{L}_{v}$ (without $\mathcal{L}_{p}$). One possible explanation for this observation is that including $\mathcal{L}_{o}$ in the loss function biases the model towards optimizing local orientation, which may affect the accurate estimation of global velocity. On the other hand, integrating $\mathcal{L}_{p}$ and $\mathcal{L}_{v}$ has a positive effect on improving the performance of NeurIT. Overall, the usage of these three loss functions strikes a balance between velocity vectors, trajectory errors, and heading orientations, allowing the model to predict the robot’s trajectory most accurately.
@@ VI-C Ablation Study
Figure 12: Analyze the impact of the data augmentation. (a) shows the performance comparison on the test set with and without data augmentation. (b) and (c) are visualization results of hidden features, which are generated by Feedforward Block, using t-SNE. We take orientation as the features and separate it into 10 bins. The visualization results demonstrate that by using data augmentation, the model can have a better generalization capability of orientation on the floor plane.
VI-D Augmentation Study
In §III-E, we discussed the implementation of a data augmentation to improve the generalization ability on all orientations of the floor plane. In this section, we present comparative experiments on data augmentation to further validate our assumption. Fig. 12a illustrates the effect of data augmentation on ATE and RTE, which demonstrates that without data augmentation, the tracking error is approximately twice as large compared to when data augmentation is applied.
We perform extra experiments to explore how well neural networks can interpret IMU data across various heading directions. The 360-degree plane is divided into 10 groups, each covering a 36-degree range. The input data is classified into ten classes based on these groups. We focus on the output of the feedforward module, before the output projection layer. Then, we utilize t-SNE [36] to visualize the reduced dimensions of the extracted data in Fig. 12b and Fig. 12c. In Fig. 12c, the boundaries between groups appear blurred or mixed, particularly between Groups 3 and 4, which indicates that the model’s ability to predict the robot’s movement direction is imprecise without data augmentation. In contrast, Fig. 12b shows much clearer boundaries between different groups.
@@ VI-D Augmentation Study
VI-E Magnetometer Study
In §II, we propose an orientation compensation method that differentiates the body-frame magnetometers as input features. It is mentioned that previous work [16, 12] used magnetometers for drift compensation as well, but in a complementary filter manner. To confirm the effectiveness of our method, we conducted two sets of comparative experiments. One uses the body-frame magnetometer as input without differentiation. The other employs a magnetometer-based complementary filter to correct the orientation error in the predicted trajectories. We experiment with three different buildings, with each trajectory lasting for more than 200m. The results are shown in Fig. 14. Obviously, our proposed method achieves optimal results in various scenarios. The results of w/o diff have a certain decrease (30%) compared to NeurIT but are still better than CF and w/o magn. Over long-distance tracking, the performance gains from the complementary filter are marginal, demonstrating its limitations in complex situations. The improvement is only 4% compared to not using magnetometers.
@@ VI-E Magnetometer Study
VI-F Comparison with Tango Phone and AVIA
We compared NeurIT with Tango phone’s visual-inertial algorithm and the FAST-LIO algorithm using a narrow FOV LiDAR (AVIA). In an environment with many feature points, NeurIT performs similarly to Tango phone, achieving an ATE of 1.08m compared to Tango phone’s 0.78m. However, in environments with limited visual feature points, NeurIT outperforms Tango phone. We did not report the results of AVIA-based FAST-LIO due to significant drift of hundreds of meters. Fig. 15b shows the estimated trajectories using various methods. Tango phone exhibits drift in corners with no visual points, while NeurIT remains consistent with the ground truth. The AVIA-based FAST-LIO algorithm is highly affected, with significant errors and severe drift after the second turn’s plain wall, making trajectory tracking impossible for the robot.
VI-G Training Datasize
We also conduct experiments on the influence of the size of the training set on the model performance, and the experimental results are shown in Fig. 16. As the size of the training set decreased from 100% to 50%, the accuracy of the model to the tracking of the robot gradually decreased, in which ATE increased from 1.08m to 1.89m, and RTE increased from 0.68m to 1.18m. The gray dashed line indicates the test results without taking magnetometers as input features. The findings reveal that when employing the magnetometers, approximately 60% of the training set attains comparable model performance as the case where the magnetometer is not implemented. Notably, even with a mere 50% of the training set, the model exhibits superior results in the realm of RTE when compared to the outcomes obtained without utilizing a magnetometer. These outcomes indicate the stability and remarkable generalization capability of our proposed methodology.
VI-H System Latency
TABLE IV: Power consumption and system complexity. The average time indicates the time consumption for predicting 1-second IMU data.
@@ VI-H System Latency
To examine NeurIT as a real-time, power-efficient solution, we evaluate its end-to-end system complexity and power consumption on NVIDIA Jetson Orin Nano, which has been widely applied in robotic applications [37, 38], in a multi-ROS system with Intel NUC. As shown in Tab. IV, NeurIT requires 34.48ms to predict 1-second IMU data in GPU mode, including 1.62ms dedicated to coordinate system transformation and noise reduction. The results promise NeurIT as a real-time, power-efficient system to run on robotic things.
VII Related Works
Inertial tracking has been proposed for many years. The Kalman filter [39], introduced in 1960, improved vehicle guidance and navigation control, including spacecraft orbit control in the Apollo program. Additional advancements, such as the extended Kalman filter [7], facilitated accurate navigation in nonlinear systems. However, when applied to low-cost IMUs on mobile phones, these algorithms perform poorly due to noise and complex pedestrian movements. To enhance phone attitude estimation, $A^{3}$ [40] utilized accelerations and magnetometers to calibrate gyroscope drift. Subsequent work, like MUSE [16], achieved promising results in device orientation tracking. These applications primarily focus on device orientation in specific areas, requiring careful calibration or parameter selection for optimal performance.
To realize indoor inertial tracking, machine learning and data-driven techniques have been exploited. RIDI [20], IONet [10], RONIN [9], and MotionTransformer [41] are examples of methods that employ deep learning and neural networks to estimate pedestrian motion trajectories. These methods have limitations in terms of generalization and applicability to other moving targets. Some studies [42] have combined neural networks with Kalman filters to improve prediction accuracy by learning noise parameters. However, the performance of existing methods still has room for optimization, particularly in handling unseen data and environments. NeurIT takes a further step in this direction and pushes the limit of the state-of-the-art accuracy.

### v2

@@ VI-B Baselines
ATE 16.75 26.04 1.72 1.76 1.72 1.66 3.94 1.63 1.40% ↑ 1.53 7.64%↑
RTE 17.58 37.51 2.04 2.09 2.17 2.16 4.34 2.04 0.16% ↑ 2.02 1.19% ↑
PDE 0.203 0.684 0.0328 0.0319 0.0345 0.0338 0.0688 0.0317 0.55% ↑ 0.0286 10.41% ↑
Seen AYE 127.97 114.15 51.39 57.52 54.71 70.95 63.74 27.21 47.05% ↑ 26.33 48.76% ↑
ATE 17.30 24.14 1.87 2.10 1.96 1.99 3.20 1.51 19.10% ↑ 1.51 19.10% ↑
RTE 18.02 40.67 1.87 2.18 1.87 1.92 3.20 1.66 11.10% ↑ 1.60 14.31% ↑
PDE 0.1799 0.7095 0.0291 0.0354 0.0297 0.0310 0.0577 0.0278 4.62% ↑ 0.0243 16.55% ↑
RIDI Unseen AYE 130.59 105.93 48.47 53.67 51.82 67.42 64.64 24.34 49.78% ↑ 24.21 50.05% ↑
ATE 22.86 33.26 7.76 6.20 6.18 5.68 8.78 4.58 19.38% ↑ 4.37 23.10% ↑
RTE 30.49 35.10 3.19 3.17 3.04 2.77 6.43 2.60 6.03% ↑ 2.59 6.53% ↑
PDE 0.124 0.107 0.0432 0.0390 0.0312 0.0267 0.0493 0.0216 18.99% ↑ 0.0181 31.99% ↑
Seen AYE 122.48 120.96 63.12 74.49 69.059 107.92 86.91 46.56 26.24% ↑ 47.60 24.59% ↑
ATE 17.30 33.47 8.91 7.73 7.57 6.38 10.27 5.78 9.52% ↑ 5.35 16.26% ↑
RTE 18.02 41.07 4.12 4.63 3.85 3.86 8.20 3.59 6.85% ↑ 3.55 7.89% ↑
PDE 0.0912 0.3166 0.0462 0.0358 0.03622 0.0320 0.0592 0.0268 16.25% ↑ 0.0243 24.06% ↑
RONIN Unseen AYE 110.00 114.80 64.46 76.49 69.21 106.80 88.69 47.88 25.72% ↑ 46.66 27.61% ↑
VI-C Evaluation Metrics
  • Absolute Trajectory Error (ATE): ATE [50] calculates the root mean square error (RMSE) of the predicted trajectory and the ground truth trajectory, which is given as $\sqrt{\frac{1}{n}\sum_{i}\|\mathbf{p}_{i}-\mathbf{p}^{g}_{i}\|_{2}^{2}}$.
  • Relative Trajectory Error (RTE): RTE [50] measures the RMSE between the predicted trajectory and the ground truth over a time interval. RTE can be calculated as $\sqrt{\frac{1}{n}\sum_{i}\left\|(\mathbf{p}_{i+\Delta t}-\mathbf{p}_{i})-(\mathbf{p}^{g}_{i+\Delta t}-\mathbf{p}^{g}_{i})\right\|_{2}^{2}}$, where $\Delta t$ is set to 60s in our evaluation.
  • Position Drift Error (PDE): PDE [14] computes the final position drift over the whole sequence. It can be expressed as $\|\mathbf{p}_{N}-\mathbf{p}^{g}_{N}\|_{2}/{D}$, where $N$ is the last index of the sequence, and $D$ denotes the ground truth trajectory length.
  • Absolute Yaw Error (AYE): AYE [24] measures the RMSE between the heading orientations between ground truth and predicted trajectories, i.e., $\sqrt{\frac{1}{n}\sum_{i}\left\|\frac{\mathbf{v}_{i}}{\|\mathbf{v}_{i}\|_{2}}-\frac{\mathbf{v}^{g}_{i}}{\|\mathbf{v}^{g}_{i}\|_{2}}\right\|_{2}^{2}}$.
Figure 7: CDF of ATE and RTE on NeurIT Test Dataset
Figure 8: Examples of Predicted Trajectories of NeurIT
VII Evaluation
VII-A Overall Performance
Evaluation on NeurIT Dataset: Tab. II summarizes the overall evaluation on the NeurIT dataset. NeurIT achieves outstanding accuracy on both seen and unseen test sets, with a PDE of 0.08, equivalent to just 0.8cm per meter traveled. It reduces the average ATE by 62% compared to the best baseline, and maintains a minimal RTE of 0.6%, corresponding to 0.6m drift over 1 minute, whereas UniTS reports over double that error ($\sim$1.5m). LLIO ranks second among baselines, but shows a performance gap likely due to less effective feature design. Thanks to multi-loss learning, NeurIT also improves AYE, reducing heading error from 43° (IONet) to 33°, demonstrating better orientation estimation. Even without magnetometer inputs, NeurIT outperforms the best baseline by 30% across four metrics. When differential magnetometers are incorporated, performance improves further, particularly RTE, which is cut by nearly half, highlighting the effectiveness of our sensor fusion design.
@@ VII-A Overall Performance
Fig. 7 shows the cumulative distribution of ATE and RTE, where NeurIT consistently outperforms all baselines, especially in unseen environments. Even in the worst-case scenario, NeurIT exceeds the best-case performance of other models. Trajectory visualizations in Fig. 8 confirm NeurIT ’s stable tracking, with predicted paths closely matching ground truth, unlike other methods that show significant deviations.
Additional comparisons reveal the limitations of using only RNNs or Transformers. While UniTS performs well overall, it suffers from high AYE and occasional failure cases, leading to long-tail drift. In contrast, NeurIT effectively combines RNN and Transformer strengths, achieving state-of-the-art accuracy and robustness.
Evaluation on RIDI & RONIN Datasets: While this work primarily focuses on robotic tracking, we also evaluate NeurIT on pedestrian tracking datasets to assess its generalization capability. As shown in Tab. III, NeurIT outperforms all baseline models on both the RIDI and RONIN datasets. Although baselines perform reasonably well on seen data, their accuracy drops significantly on unseen test sets, where NeurIT achieves notably better results. On RIDI, the performance gain rises from 17% on the seen set to nearly 25% on the unseen set. For the RONIN dataset, only half of the total data is publicly available, limiting the model’s ability to learn comprehensive pedestrian features. This constraint leads to a smaller improvement in unseen conditions and generally weaker performance compared to RIDI. Overall, the strong results across both domains highlight NeurIT ’s potential as a unified model for both robotic and pedestrian inertial tracking.
VII-B Benchmark Study
Performance over Time and Distance: We analyze how prediction errors grow over time and distance. As shown in Fig. 9a and Fig. 9b, error increases with longer tracking durations and distances, as expected. However, NeurIT shows a significantly slower error growth compared to the best baseline, UniTS. For instance, after 250 meters, NeurIT’s ATE remains lower than UniTS’s ATE at just 50 meters. Over a 5-minute or 250-meter trajectory, NeurIT maintains a mean error of about 0.9m, whereas UniTS reaches the same error within 1 minute and exceeds 2.6m after 5 minutes. These results demonstrate the strong stability and robustness of NeurIT for long-duration indoor tracking.
@@ VII-B Benchmark Study
Velocity Prediction Error: To evaluate NeurIT ’s performance in speed estimation, we compare its predicted speed against ground truth. As shown in Fig. 10a, NeurIT consistently outperforms the best baseline (UniTS), with an average error margin of about 0.2m/s. The error remains low and stable, under 0.4m/s, until speeds exceed 1.2m/s, beyond which it grows slightly, likely due to limited high-speed samples in the training data. This suggests that including more high-speed instances could further improve accuracy. In contrast, UniTS performs poorly across all speeds, with an average error exceeding 2m/s. As illustrated in Fig. 10b, NeurIT generates a smooth and accurate trajectory that closely follows the ground truth, while the baseline produces erratic and imprecise movements.
Different Motion Status: Robotic movement often adapts to environmental conditions. To assess NeurIT ’s robustness across varied indoor settings, we conduct additional experiments and visualize the predicted trajectories and corresponding speeds in Fig. 11. In Building A, floor 1, a narrow, crowded office space, NeurIT accurately tracks the robot through tight corridors, including turns and straight segments. It also handles challenging cases like sudden starts and high-speed linear motion with ease. Building A, floor 5, presents a more extreme scenario, featuring a 40-meter-long narrow corridor. Even in new environments such as Buildings B and C, representing complex and open spaces, NeurIT maintains high accuracy despite not having seen these conditions during training. These results demonstrate NeurIT ’s strong robustness and adaptability across diverse and unseen indoor environments.
@@
Figure 13: Analyze the impact of the data augmentation. (a) shows the performance comparison on the test set with and without data augmentation. (b) and (c) are visualization results of hidden features generated by the Feedforward Block, using t-SNE. We take orientation as the feature and separate it into 10 bins. The visualization results demonstrate that by using data augmentation, the model can have a better generalization capability of orientation on the floor plane.
VII-C Ablation Study
We evaluate the effectiveness of key design in NeurIT, focusing on four aspects: time-frequency input (TF), block-recurrent attention module (BR), multi-loss learning (m-loss), and a detailed analysis of the impact of each loss function.
@@ VII-C Ablation Study
Figure 15: Examples of Predicted Trajectories in Two Large Buildings.
TABLE IV: Extended evaluation in large buildings.
Metric UniTS
NeurIT
w/o magn
Impvt NeurIT Impvt
ATE 2.70 1.66 38.41% ↑ 1.36 49.79% ↑
RTE 2.27 1.49 34.65% ↑ 1.18 48.10% ↑
@@
TABLE V: Model Complexity.
Model
# Params
(10^6)
Avg. GPU
time (ms)
GFLOPs
(10^9/s)
NeurIT 1.9 6.09 0.6
IONet 0.11 2.16 0.109
RONIN 0.22 4.04 0.107
UniTS 165.9 16.93 19.6
TABLE VI: Power consumption and system complexity.
@@
Multi-loss Learning: To evaluate the effectiveness of multi-loss learning, we first replace it with a single position loss, as used in RONIN. This leads to a noticeable performance drop across all four metrics. We observe some counterintuitive results: adding the orientation loss $\mathcal{L}_{\mathbf{o}}$ to the position loss $\mathcal{L}_{\mathbf{p}}$ improves heading accuracy (AYE decreases from 35° to 30°) but degrades ATE and RTE. A similar trend is seen when combining $\mathcal{L}_{\mathbf{o}}$ with velocity loss $\mathcal{L}_{\mathbf{v}}$ while excluding $\mathcal{L}_{\mathbf{p}}$. This suggests that emphasizing orientation loss may bias the model toward local rotation accuracy at the expense of global trajectory estimation. In contrast, combining $\mathcal{L}_{\mathbf{p}}$ and $\mathcal{L}_{\mathbf{v}}$ improves overall performance. The full multi-loss setting strikes a balance between velocity, trajectory, and orientation, leading to the most accurate trajectory predictions.
VII-D Augmentation Study
As discussed in §IV-E, we apply data augmentation to improve generalization across all floor-plane orientations. To validate its effectiveness, we conduct comparative experiments. Fig. 13a shows that without augmentation, both ATE and RTE nearly double compared to the augmented setting.
To further analyze how well the model captures heading direction, we divide the 360-degree plane into 10 groups (each spanning 36°) and label the input data accordingly. We extract the hidden embeddings from the feedforward module (before output projection) and visualize them using t-SNE [51] in Fig. 13b and Fig. 13c. Without augmentation, group boundaries are blurred, particularly between Groups 3 and 4, indicating poor directional discrimination. In contrast, Fig. 13b shows clear separation between groups, confirming that data augmentation significantly improves the model’s ability to interpret orientation-dependent IMU signals.
VII-E Magnetometer Study
In NeurIT, we propose using differentiation of body-frame magnetometers for orientation compensation. To validate this component, we compare three approaches: using magnetometers without differentiation, applying a magnetometer-based complementary filter (CF) for drift compensation [16, 12], and using our method. Tests on 200m+ trajectories across three buildings show our method achieves the best results (Fig. 14). The undifferentiated approach sees a 30% drop, while the CF approach offers only a 4% improvement over not using magnetometers, rendering its limitations in complex scenarios.
VII-F Extended Experiments in Large Urban Complexes
To further evaluate NeurIT in complex real-world settings, we conduct extended tests in two large urban environments: a shopping mall and a concert hall. We collect 10 samples per location, each lasting approximately 5 minutes, and compare NeurIT with UniTS, the strongest baseline. As shown in Tab. IV, NeurIT consistently outperforms UniTS on both ATE and RTE metrics, with nearly a 50% improvement. Our magnetometer-based compensation also yields an average accuracy gain of 15% over models without magnetometer input. Trajectory visualizations in Fig. 15 highlight UniTS’s drift over time, while NeurIT maintains stable and accurate tracking across diverse indoor environments.
VII-G Comparison with Tango Phone and AVIA
To evaluate how NeurIT compares with established visual-inertial and LiDAR-based tracking systems, we tested it against the Tango phone’s visual-inertial algorithm and the FAST-LIO algorithm using a narrow FOV LiDAR (AVIA). In environments rich with visual features, NeurIT achieves comparable performance to the Tango phone, with an ATE of 1.08m versus 0.78m. However, in visually sparse settings, NeurIT outperforms the Tango phone, which suffers from drift in featureless corners. Fig. 16b illustrates the trajectory comparisons, where NeurIT aligns with the ground truth.
VII-H Training Datasize
We further examine how training set size affects model performance, as shown in Fig. 17. Reducing the training set from 100% to 50% leads to a drop in accuracy, with ATE increasing from 1.08m to 1.89m and RTE from 0.68m to 1.18m. Interestingly, using magnetometer data allows the model to match the performance of the non-magnetometer variant with only 60% of the data. Even at 50% data, it still achieves better RTE. These results demonstrate the robustness and strong generalization of our approach.
VII-I System Latency
Model Complexity: We evaluate the complexity of selected models on an NVIDIA 4090 GPU. Tab. V reports each model’s parameter count, average GPU execution time for 1-second data, and GFLOPs during testing. NeurIT not only delivers the best performance but is also one of the smallest models—significantly smaller than the best baseline, UniTS. This demonstrates that NeurIT strikes a good balance between model size and computational efficiency.
@@ VII-I System Latency
System Complexity: To examine NeurIT as a real-time, power-efficient solution, we evaluate its end-to-end system complexity and power consumption on NVIDIA Jetson Orin Nano, which has been widely applied in robotic applications [52, 53], in a multi-ROS system with Intel NUC. As shown in Tab. VI, NeurIT requires 34.48ms to predict 1-second IMU data in GPU mode, including 1.62ms dedicated to coordinate system transformation and noise reduction. The results promise NeurIT as a real-time, power-efficient system to run on robotic things.
VII-J Limitations
While NeurIT consistently outperforms baseline models, our evaluation reveals two key limitations. First, at high speeds (>1.2,m/s), velocity prediction errors increase slightly, likely due to limited high-speed samples and increased IMU noise from wheel vibrations. Second, long-duration trajectories show gradual drift caused by small cumulative velocity errors. Future work could address these issues by augmenting high-speed training data, applying adaptive filtering to reduce noise, and incorporating loop-closure mechanisms to correct long-range drift in revisited indoor locations.